# MoH: Multi-Head Attention as Mixture-of-Head Attention

### v1

@@ 4.2 DiT for Class-Conditional Image Generation
Methods #Activated Language Tasks Avg.
 Heads (%) SciQ PIQA WinoGrande OpenbookQA LogiQA TruthfulQA
@@
4.3 Training LLMs from Scratch
Model Settings.    For training LLMs from scratch, we use Megatron (Shoeybi et al., 2019), an open-source training code, as the training framework. Please refer to the Appendix for detailed hyper-parameter settings (Tab. A) of various MoH-LLMs. All models are trained with the AdamW optimizer (Loshchilov & Hutter, 2017), using a batch size of 4 million tokens with a sequence length of 2048. The final learning rate is set to 10% of the maximum. During training, a weight decay of 0.1 and gradient clipping of 1.0 are applied. For LLM-S and MoH-LLM-S, the maximum learning rate is set to 3e-4. For LLM-B and MoH-LLM-B, the maximum learning rate is set to 5e-4.
Training Details.    We only use public datasets for training, ensuring accessibility for academic research. Specifically, we sample from the RedPajama (Computer, 2023), Dolma (Soldaini et al., 2024), and Pile (Gao et al., 2020) datasets according to different sampling probabilities. Please refer to the Appendix for detailed sample ratios (Tab. B). Following previous works, we utilize the tokenizer from LLaMA2 (Touvron et al., 2023), which contains 65,536 vocabulary tokens.
Evaluation Benchmarks.    The evaluation is performed on multiple benchmarks using the Eleuther AI Language Model Evaluation Harness (Gao et al., 2024), a unified framework for testing generative language models. Since the parameters are only about 0.2B for the smallest model, we select 6 simple benchmarks as the metric. Specifically, we report 0-shot accuracy on SciQ (Welbl et al., 2017), PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2021), OpenbookQA (Mihaylov et al., 2018), LogiQA (Liu et al., 2020), and TruthfulQA (Lin et al., 2022).
Table 5: Comparisons between MoH-LLaMA3-8B and LLaMA3-8B. Please refer to the Appendix for the performance of the model at the end of the first stage of training.
Methods #Activated MMLU (5) CEVAL (5) CMMLU (5) GSM8K(8) TruthfulQA
 Heads (%)
@@ 4.3 Training LLMs from Scratch
Methods #Activated HellaSwag (10) LogiQA BoolQ (32) LAMBADA SciQ
 Heads (%)
@@
Methods #Activated PIQA WinoGrande NQ (32) ARC-C (25) Average
 Heads (%)
@@
4.4 Continue-Tuning LLaMA3-8B
$g_{i}^{q}=\mathds{1}(\text{Token $\bm{x}$ selects Head $i$}),$ (9)
where $\mathds{1}(*)$ denotes the indicator function. $g_{i}^{q}$ represents the quantized routing score. We then adopt a straight-through estimator, which assigns the incoming gradients to a threshold operation to be the outgoing gradients, which is formulated as:
where $g_{i}$ denotes the real-valued routing score. This simple approximation function significantly mitigates the issue of gradient vanishing (Wang et al., 2024). Similar to training LLMs from scratch, we also use Megatron (Shoeybi et al., 2019), an open-source training code, as the training framework.
Training Details.    We find that if there is a discrepancy between the continue-training data and the original training data distribution of the model, the performance of the model may fluctuate wildly at the beginning of the training process. Since we are unable to have access to the raw training data of LLaMA3, we address these potential performance fluctuations by dividing the training process into two stages. In the first stage, we continue-tune the original LLaMA3-8B model using 300B tokens to adapt the model to our dataset. In the second stage, we continue-tune this adapted model into our proposed MoH model with 100B tokens. During the first stage, the maximum learning rate is set to 6e-5, and the final learning rate is 6e-6. In the second stage, the maximum learning rate is set to 2e-5, and the final learning rate is 1e-6. For both stages, we employ the AdamW optimizer (Loshchilov & Hutter, 2017), with a batch size of 16 million tokens with a sequence length of 8192. During training, we use a weight decay of 0.1 and gradient clipping of 1.0.
Evaluation Benchmarks.    We use the Eleuther AI Language Model Evaluation Harness (Gao et al., 2024) to evaluate models on multiple key benchmarks. Specifically, we utilize the lm-evaluation-harness package to assess performance on a comprehensive suite of downstream tasks: (i) Following Pythia (Biderman et al., 2023), we report 0-shot accuracy on LAMBADA (Paperno et al., 2016), LogiQA (Liu et al., 2020), PIQA (Bisk et al., 2020), SciQ (Welbl et al., 2017), and WinoGrande (Sakaguchi et al., 2021). (ii) We report the accuracy of Chinese tasks, including 5-shot CEVAL (Huang et al., 2023) and 5-shot CMMLU (Li et al., 2023a). (iii) We report the accuracy of tasks from the Open LLM Leaderboard (Beeching et al., 2023), including 10-shot HellaSwag (Zellers et al., 2019), 25-shot ARC Challenge (ARC-C) (Clark et al., 2018), and 5-shot MMLU (Hendrycks et al., 2021). (iv) We report the exact match score for 32-shot Natural Questions (NQ) (Kwiatkowski et al., 2019) and the accuracy for 32-shot BoolQ (Clark et al., 2019). (v) We report the exact match score for 8-shot GSM8K (Cobbe et al., 2021) to evaluate the math ability. (vi) Moreover, we report 0-shot accuracy on TruthfulQA (Lin et al., 2022) to assess the ability to generate truthful answers.
@@ 4.5 Ablative Analysis
Figure 3: Visualization of the head load distribution in the final MoH layer. For ViT and DiT, we present the head load distributions for the categories “Desk”, “Goldfish”, and “Ice cream”. For LLM, we display the head distributions for the tasks “LogiQA”, “PIQA”, and “WinoGrande”. MoH-ViT-B, MoH-DiT-XL/2, and MoH-LLM-B activate 75%, 90%, and 75% of the attention heads, respectively.
@@ 5 Discussion
The Difference between MoH and MoA.    We clarify the differences between MoH and MoA (Zhang et al., 2022) from the following three aspects. First, in terms of motivation, the goal of MoH is to improve the efficiency and performance of the attention mechanism without increasing the number of parameters. In contrast, MoA shares the motivation of MoE, which is to expand model parameters while keeping inference costs low. Therefore, the model settings of MoH are more stringent than those of MoA. Second, in terms of methodology, our MoH introduces shared heads and two-stage routing to enhance the standard MoE method. More importantly, we show that pre-trained multi-head attention models can be further continue-tuned into our MoH models, greatly improving the applicability of the proposed MoH method. In contrast, MoA directly combines multi-head attention with MoE. Due to the adoption of shared keys and values, MoA must be trained from scratch, which limits its applicability. Finally, in terms of model frameworks, our MoH is validated across various popular model frameworks and tasks, including ViT, DiT, and decoder-only LLMs, while MoA is only validated on the encoder-decoder architecture for language tasks.
@@ 6 Conclusion
In this paper, we introduce MoH, a promising alternative to multi-head attention. MoH enables each token to adaptively select the appropriate attention heads, improving both model performance and inference efficiency without increasing the number of parameters. Extensive experiments across various popular model frameworks, including ViT, DiT, and LLMs, demonstrate that MoH outperforms multi-head attention, even when using only 50$\sim$90of the attention heads. More encouragingly, we show that pre-trained multi-head attention models, such as LLaMA3-8B, can be further continue-tuned into our MoH models, significantly enhancing the applicability of the proposed MoH method. This work represents a promising step toward advanced and efficient attention-based models, which may be meaningful and helpful to both the research and industrial communities.
Appendix A Appendix
Abstract.    This appendix provides additional discussions (Appendix A), implementation details (Appendix B), several additional experiments (Appendix C), more qualitative analysis (Appendix D), and details of quantitative evaluations for LLMs (Appendix E).
Code.    We have attached the code to the supplementary material. In this code, we also provide the evaluation process of the proposed method. We promise to release a more detailed and clean code version upon publication.
@@ Appendix A Additional Discussions
A.1 Limitations and Future Work
@@
Multimodal Inputs.    Effectively processing information from multiple modalities in the attention mechanism remains an open question. Recent work (Wan et al., 2024) has shown that visual and textual tokens exhibit distinct attention patterns in multi-head attention. Future work could explore the attention patterns of MoH with different modal inputs, for example within multimodal large language models (Jin et al., 2024a; Lin et al., 2023; 2024; Liu et al., 2024; Jin et al., 2023b).
More Downstream Tasks.    We evaluate our proposed MoH across various popular model frameworks, including ViT for image classification, DiT for class-conditional image generation, and LLMs for language tasks. Future work can explore the application of MoH in more downstream tasks, such as audio tasks and multimodal tasks (Jin et al., 2023a; 2022).
@@ Appendix B Implementation Details
B.1 Training LLMs from Scratch
Model Settings.    For training LLMs from scratch, we use Megatron (Shoeybi et al., 2019), an open-source training code, as the training framework. The detailed hyper-parameter settings of various MoH-LLMs are shown in Tab. A.
Table A: Sizes and architectures of MoH-LLMs and LLMs. “MoH-LLM-B” has more parameters than “LLM-B” due to the additional parameters introduced by the router network.
@@ B.1 Training LLMs from Scratch
Data Details.    Consistent with previous works, we use the tokenizer of LLaMA2, which contains 65,536 vocabulary tokens. It is worth noting that MoH-LLM is trained exclusively on public datasets, making it accessible for academic research settings. Tab. B shows the detailed sample ratios of different open-source datasets. Specifically, we sample from the following datasets according to different sampling probabilities:
@@
Table B: Sampling ratio of different open-source datasets for MoH-LLMs. MoH-LLM is trained exclusively on public datasets, making it accessible for academic research settings.
@@
Training Hyper-Parameters.    Tab. C shows the detailed training hyper-parameters of MoH-LLMs. Specifically, all MoH-LLMs are trained with the AdamW optimizer (Loshchilov & Hutter, 2017), using a batch size of 4 million tokens with a sequence length of 2048. The final learning rate is set to 10% of the maximum. During training, a weight decay of 0.1 and gradient clipping of 1.0 are applied. For LLM-S and MoH-LLM-S, the maximum learning rate is set to 3e-4. For LLM-B and MoH-LLM-B, the maximum learning rate is set to 5e-4.
Table C: Training hyper-parameters of MoH-LLMs.
@@
B.2 Continue-Tuning LLaMA3-8B
Training Hyper-Parameters.    Tab. D shows the detailed training hyper-parameters of MoH-LLaMA3-8B. We find that if there is a discrepancy between the continue-training data and the original training data distribution of the model, the performance of the model may fluctuate wildly at the beginning of the training process. Since we do not have access to the raw training data of LLaMA3, we address these potential performance fluctuations by dividing the training process into two stages. In the first stage, we continue-tune the original LLaMA3-8B model using 300B tokens to adapt it to our dataset. In addition, during the first stage, to enhance the Chinese ability of the model, we expand the vocabulary size. Specifically, we increase the original LLaMA3-8B vocabulary size from 128,256 to 160,896. In the second stage, we continue-tune this adapted model into our proposed MoH model with 100B tokens. During the first stage, the maximum learning rate is set to 6e-5, and the final learning rate is 6e-6. In the second stage, the maximum learning rate is set to 2e-5, and the final learning rate is 1e-6. For both stages, we employ the AdamW optimizer (Loshchilov & Hutter, 2017), with a batch size of 16 million tokens with a sequence length of 8192. During training, we use a weight decay of 0.1 and gradient clipping of 1.0.
Table D: Training hyper-parameters of MoH-LLaMA3-8B. We divide the training process into two stages. In the first stage, we continue-tune the LLaMA3-8B model using 300B tokens. In the second stage, we continue-tune this adapted model into our proposed MoH model with 100B tokens.

### v2

@@ 4.1 ViT for Image Classification
Methods #Activated Heads (%) SciQ PIQA WinoGrande OpenbookQA LogiQA TruthfulQA Average
@@
Table 4: Comparisons between MoH-LLaMA3-8B and LLaMA3-8B. Please refer to Tab. G in the Appendix for the performance of the model at the end of the first stage of training.
Methods #Activated Heads (%) MMLU (5) CEVAL (5) CMMLU (5) GSM8K(8) TruthfulQA
@@
Methods #Activated Heads (%) MMLU (5) CEVAL (5) CMMLU (5) GSM8K(8) TruthfulQA
@@
Methods #Activated Heads (%) MMLU (5) CEVAL (5) CMMLU (5) GSM8K(8) TruthfulQA
@@
Results.    As shown in Tab. 1, despite activating only a subset of attention heads, MoH-ViT achieves highly competitive performance compared to current state-of-the-art methods. For example, MoH-ViT-B achieves 84.9% Top-1 accuracy on the ImageNet-1K classification benchmark with just 75% of the attention head. In contrast, the well-established ViT baseline, TransNeXt, attains a slightly lower accuracy of 84.8% while requiring 100% of the heads to be activated. These results suggest that MoH is a promising alternative to multi-head attention for vision model design.
@@
Table 7: Comparisons about inference time. We convert the $\bm{Q}$, $\bm{K}$, and $\bm{V}$ features into sparse matrices using the mask generated by the router and replace the dense matrix multiplication in the attention mechanism with sparse matrix multiplication. To eliminate the impact of underlying operator optimizations, we replaced all matrix multiplications with sparse matrix multiplication when testing for speed.
Methods #Head Num #Head Dim #Sequence Length #Activated Heads (%) Time (ms)
Multi-Head Attention 32 64 256 100 0.360
MoH (Ours) 32 64 256 90 0.352
MoH (Ours) 32 64 256 75 0.321
MoH (Ours) 32 64 256 50 0.225
Multi-Head Attention 32 64 512 100 1.376
MoH (Ours) 32 64 512 90 1.351
MoH (Ours) 32 64 512 75 1.180
MoH (Ours) 32 64 512 50 0.863
4.2 DiT for Class-Conditional Image Generation
Model Settings.    For Diffusion models with Transformers (DiT) (Peebles & Xie, 2023), we only replace the standard multi-head attention with our MoH in MoH-DiT models, while keeping all other training parameters identical to DiT. We use the ImageNet-1K dataset for class-conditional image generation at a resolution of 256$\times$256. To evaluate generation performance, we use Frechet Inception Distance (FID(Heusel et al., 2017) to assess overall sample quality, Precision and Recall (Kynkäänniemi et al., 2019) to measure fidelity and diversity separately, and sFID (Nash et al., 2021) as a metric that better captures spatial relationships than FID. Moreover, we use Inception Score (IS(Salimans et al., 2016) as another metric for fidelity.
Results.    As shown in Tab. 2, MoH-DiT models consistently outperform DiT models with 90% of heads activated. However, when only 75% of the heads are activated, MoH-DiT models perform worse than DiT models. This may be because image generation tasks are dense prediction tasks that require attention mechanisms to capture pixel-level fine-grained relationships, leaving less redundancy in the attention heads compared to image classification tasks. These results suggest that MoH is a promising alternative to multi-head attention for diffusion models.
4.3 Training LLMs from Scratch
Model Settings.    For training LLMs from scratch, we use Megatron (Shoeybi et al., 2019), an open-source training code, as the training framework. Please refer to the Appendix for detailed hyper-parameter settings (Tab. C) of various MoH-LLMs. The evaluation is performed on multiple benchmarks using the Eleuther AI Language Model Evaluation Harness (Gao et al., 2024), a unified framework for testing generative language models. Since the parameters are only about 0.2B for the smallest model, we select 6 simple benchmarks as the metric. Specifically, we report 0-shot accuracy on SciQ (Welbl et al., 2017), PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2021), OpenbookQA (Mihaylov et al., 2018), LogiQA (Liu et al., 2020), and TruthfulQA (Lin et al., 2022).
Training Details.    We only use public datasets for training, ensuring accessibility for academic research. Specifically, we sample from the RedPajama (Computer, 2023), Dolma (Soldaini et al., 2024), and Pile (Gao et al., 2020) datasets according to different sampling probabilities. Please refer to the Appendix for detailed sample ratios. Following previous works (Jin et al., 2025), we utilize the tokenizer from LLaMA2 (Touvron et al., 2023), which contains 65,536 vocabulary tokens.
4.4 Continue-Tuning LLaMA3-8B
$g_{i}^{q}=\mathds{1}(\text{Token $\bm{x}$ selects Head $i$}),$ (9)
where $\mathds{1}(*)$ denotes the indicator function. $g_{i}^{q}$ represents the quantized routing score. We then adopt a straight-through estimator, which assigns the incoming gradients to a threshold operation to be the outgoing gradients:
where $g_{i}$ denotes the real-valued routing score. This approximation function significantly mitigates the issue of gradient vanishing (Wang et al., 2024). Similar to training LLMs from scratch, we also use Megatron (Shoeybi et al., 2019), an open-source training code, as the training framework.
Figure 3: Visualization of the head load distribution in the final MoH layer. For ViT and DiT, we present the head load distributions for the categories “Desk”, “Goldfish”, and “Ice cream”. For LLM, we display the head distributions for the tasks “LogiQA”, “PIQA”, and “WinoGrande”. MoH-ViT-B, MoH-DiT-XL/2, and MoH-LLM-B activate 75%, 90%, and 75% of the attention heads, respectively. “Density” denotes the ratio of the number of head activations to the total number of tokens.
Training Details.    We find that if there is a discrepancy between the continue-training data and the original training data distribution of the model, the performance of the model may fluctuate wildly at the beginning of the training process. Since we are unable to have access to the raw training data of LLaMA3, we address these potential performance fluctuations by dividing the training process into two stages. In the first stage, we continue-tune the original LLaMA3-8B model using 300B tokens to adapt the model to our dataset. In the second stage, we continue-tune this adapted model into our proposed MoH model with 100B tokens. We utilize the lm-evaluation-harness package to assess performance on a comprehensive suite of downstream tasks: (i) Following Pythia (Biderman et al., 2023), we report 0-shot accuracy on LAMBADA (Paperno et al., 2016), LogiQA (Liu et al., 2020), PIQA (Bisk et al., 2020), SciQ (Welbl et al., 2017), and WinoGrande (Sakaguchi et al., 2021). (ii) We report the accuracy of Chinese tasks, including 5-shot CEVAL (Huang et al., 2023) and 5-shot CMMLU (Li et al., 2023a). (iii) We report the accuracy of tasks from the Open LLM Leaderboard (Beeching et al., 2023), including 10-shot HellaSwag (Zellers et al., 2019), 25-shot ARC Challenge (ARC-C) (Clark et al., 2018), and 5-shot MMLU (Hendrycks et al., 2021). (iv) We report the exact match score for 32-shot Natural Questions (NQ) (Kwiatkowski et al., 2019) and the accuracy for 32-shot BoolQ (Clark et al., 2019). (v) We report the exact match score for 8-shot GSM8K (Cobbe et al., 2021) to evaluate the math ability. (vi) Moreover, we report 0-shot accuracy on TruthfulQA (Lin et al., 2022) to assess the ability to generate truthful answers.
@@ 5 Discussion
The Efficiency of Our Proposed MoH.    To explore if our method performs better with longer sequences, we increase the input sequence length. For rows 1 to 4 of Tab. 7, the input length is 256. For rows 5 to 8, it is 512. As shown in Tab. 7, although dynamic routing introduces additional computational overhead, MoH still outperforms standard multi-head attention mechanisms. Furthermore, as the input sequence gets longer, the advantage of MoH grows.
@@
The Difference between MoH and MoA.    We clarify the differences between MoH and MoA (Zhang et al., 2022) from the following three aspects. First, in terms of motivation, the goal of MoH is to improve the efficiency and performance of the attention mechanism without increasing the number of parameters. In contrast, MoA shares the motivation of MoE, which is to expand model parameters while keeping inference costs low. Therefore, the model settings of MoH are more stringent than those of MoA. Second, in terms of methodology, our MoH introduces shared heads and two-stage routing to enhance the standard MoE method. More importantly, we show that pre-trained multi-head attention models can be further continue-tuned into our MoH models, greatly improving the applicability of the proposed MoH method. In contrast, MoA directly combines multi-head attention with MoE. Due to the adoption of shared keys and values, MoA must be trained from scratch, which limits its applicability. Finally, in terms of model frameworks, our MoH is validated across various popular model frameworks and tasks, including ViT, DiT, and decoder-only LLMs, while MoA is only validated for language tasks.
@@ 6 Conclusion
In this work, we introduce MoH, a promising alternative to multi-head attention. MoH enables each token to adaptively select the appropriate attention heads, improving both model performance and inference efficiency without increasing the number of parameters. Extensive experiments across various popular model frameworks, including ViT, DiT, and LLMs, demonstrate that MoH outperforms multi-head attention, even when using only 50$\sim$90of the attention heads. This work represents a promising step toward advanced and efficient attention-based models, which may be helpful to both the research and industrial communities.
Acknowledgements
This work was supported in part by the Natural Science Foundation of China (No. 62202014, 62332002, 62425101, 62088102), and NUS Start-up Grant A-0010106-00-00. Besides, this work was performed when Peng Jin was an Intern at Skywork AI.
Impact Statement
This work is an important step toward creating more advanced and efficient attention-based models, which could benefit both the research and industrial communities. Efficient attention models will not only lower the training costs for researchers but also greatly reduce the expenses involved in deploying and using large models.
@@ Appendix A Additional Discussions
A.1 Why is MoH Superior to Vanilla Multi-Head Attention?
We demonstrate that MoH is superior to vanilla multi-head attention from both theoretical and experimental perspectives.
Specifically, MoH not only improves efficiency and model performance but also helps different attention heads to specialize better compared to multi-head attention.
From the theoretical perspective, in standard multi-head attention, all heads use the same data, which can cause them to learn similar features. Many studies have pointed out that there are redundant heads in multi-head attention. Given a minibatch of data $D$, the gradient of each attention head in multi-head attention can be written as $\mathbb{E}_{x\in D}[\frac{\partial\mathcal{L}(x)}{\partial h_{i}}]$.
In contrast, in MoH, routed heads are trained only on smaller subsets of data specifically assigned to them. In MoH’s routing mechanism, the data is divided into $h-h_{s}$ subsets $\{D_{1},D_{2},...,D_{h-h_{s}}\}$, with each subset corresponding to a routed head. Besides, the routing score for each attention head acts as an adaptive adjustment to the learning rate, enabling the attention heads in MoH to specialize more effectively. Given a minibatch of data $D$ and the router $G(*)$, the gradient of each routed head in MoH can be written as $\mathbb{E}_{x\in D_{i}}[G(x)_{i}\frac{\partial\mathcal{L}(x)}{\partial h_{i}}]$. The gradient of each shared head in MoH can be written as $\mathbb{E}_{x\in D}[G(x)_{i}\frac{\partial\mathcal{L}(x)}{\partial h_{i}}]$. As shown in Tab. A, the routing mechanism and adaptive weights in MoH enable attention heads to specialize more effectively compared to standard multi-head attention.
Table A: Comparisons between the multi-head attention and our proposed mixture-of-head attention.
Methods #Head Type #Data #Weight (learning rate) #Gradient
Multi-Head Attention - $D$ 1 $\mathbb{E}_{x\in D}[\frac{\partial\mathcal{L}(x)}{\partial h_{i}}]$
MoH routed head $D_{i}\in D$ $G(x)_{i}$ $\mathbb{E}_{x\in D_{i}}[G(x)_{i}\frac{\partial\mathcal{L}(x)}{\partial h_{i}}]$
MoH shared head $D$ $G(x)_{i}$ $\mathbb{E}_{x\in D}[G(x)_{i}\frac{\partial\mathcal{L}(x)}{\partial h_{i}}]$
From the experimental perspective, we calculated the similarity of attention patterns and output features of different attention heads (include routed heads and shared heads). As shown in Tab. B, the similarity of attention patterns and output features among attention heads in MoH is lower than in standard multi-head attention, indicating reduced redundancy and greater differentiation among the attention heads in MoH.
Table B: The similarity of attention patterns and output features among attention heads. Given a pair of attention score matrices $A$ and $A^{\prime}$, we calculate the similarity of attention patterns as $1-\frac{1}{2}\mathbb{E}[||A-A^{\prime}||_{1}]$. Since attention scores form a probability distribution for each query, the similarity is always between 0 to 1.
Methods Similarity of Attention Patterns Cosine Similarity of Output Features
ViT LLM ViT LLM
Multi-Head Attention 0.5159 0.4795 0.0411 0.2550
MoH 0.3978 0.4333 0.0165 0.2042
A.2 Limitations and Future Work
@@
Multimodal Inputs.    Effectively processing information from multiple modalities in the attention mechanism remains an open question. Recent work (Wan et al., 2024) has shown that visual and textual tokens exhibit distinct attention patterns in multi-head attention. Future work could explore the attention patterns of MoH with different modal inputs, for example within multimodal large language models (Jin et al., 2024b; Lin et al., 2023, 2024; Liu et al., 2024; Jin et al., 2023, 2024a).
More Downstream Tasks.    We evaluate our proposed MoH across various popular model frameworks, including ViT for image classification, DiT for class-conditional image generation, and LLMs for language tasks. Future work can explore the application of MoH in more downstream tasks, such as audio tasks and multimodal tasks.
@@ Appendix B Implementation Details
B.1 ViT for Image Classification
B.2 DiT for Class-Conditional Image Generation
B.3 Training LLMs from Scratch
Model Settings.    For training LLMs from scratch, we use Megatron (Shoeybi et al., 2019), an open-source training code, as the training framework. The detailed hyper-parameter settings of various MoH-LLMs are shown in Tab. C.
Table C: Sizes and architectures of MoH-LLMs and LLMs. “MoH-LLM-B” has more parameters than “LLM-B” due to the additional parameters introduced by the router network.
@@ B.3 Training LLMs from Scratch
Data Details.    Consistent with previous works, we use the tokenizer of LLaMA2, which contains 65,536 vocabulary tokens. It is worth noting that MoH-LLM is trained exclusively on public datasets, making it accessible for academic research settings. Tab. D shows the detailed sample ratios of different open-source datasets. Specifically, we sample from the following datasets according to different sampling probabilities:
@@
Table D: Sampling ratio of different open-source datasets for MoH-LLMs. MoH-LLM is trained exclusively on public datasets, making it accessible for academic research settings.
@@
Training Hyper-Parameters.    Tab. E shows the detailed training hyper-parameters of MoH-LLMs. Specifically, all MoH-LLMs are trained with the AdamW optimizer (Loshchilov & Hutter, 2017), using a batch size of 4 million tokens with a sequence length of 2048. The final learning rate is set to 10% of the maximum. During training, a weight decay of 0.1 and gradient clipping of 1.0 are applied. For LLM-S and MoH-LLM-S, the maximum learning rate is set to 3e-4. For LLM-B and MoH-LLM-B, the maximum learning rate is set to 5e-4.
Table E: Training hyper-parameters of MoH-LLMs.
@@
B.4 Continue-Tuning LLaMA3-8B
Training Hyper-Parameters.    Tab. F shows the detailed training hyper-parameters of MoH-LLaMA3-8B. We find that if there is a discrepancy between the continue-training data and the original training data distribution of the model, the performance of the model may fluctuate wildly at the beginning of the training process. Since we do not have access to the raw training data of LLaMA3, we address these potential performance fluctuations by dividing the training process into two stages. In the first stage, we continue-tune the original LLaMA3-8B model using 300B tokens to adapt it to our dataset. In addition, during the first stage, to enhance the Chinese ability of the model, we expand the vocabulary size. Specifically, we increase the original LLaMA3-8B vocabulary size from 128,256 to 160,896. In the second stage, we continue-tune this adapted model into our proposed MoH model with 100B tokens. During the first stage, the maximum learning rate is set to 6e-5, and the final learning rate is 6e-6. In the second stage, the maximum learning rate is set to 2e-5, and the final learning rate is 1e-6. For both stages, we employ the AdamW optimizer (Loshchilov & Hutter, 2017), with a batch size of 16 million tokens with a sequence length of 8192. During training, we use a weight decay of 0.1 and gradient clipping of 1.0.
Table F: Training hyper-parameters of MoH-LLaMA3-8B. We divide the training process into two stages. In the first stage, we continue-tune the LLaMA3-8B model using 300B tokens. In the second stage, we continue-tune this adapted model into our proposed MoH model with 100B tokens.